OpenStudy (anonymous):

MEDAL AND FAN!! Find the x and y intercepts of the line with the given equations. a) x-y=4 b) 2x-y=10

OpenStudy (eric_d):

To find x intercept, put y=o into the equation

OpenStudy (anonymous):

solve for the a in terms of x, x=4+y substitute in b for x. giving you 2(4+y)-y=10 8+2y-y=10 8+y=10 y=2 substitute in x=4+y= 4+2=6 so x=6 y=2

OpenStudy (anonymous):

were they two different problems or one just given two equations?

OpenStudy (eric_d):

Two different equations

OpenStudy (anonymous):

oh they were 2 different equations and the directions are the same! :D

OpenStudy (anonymous):

oh do u know how to do b?

OpenStudy (eric_d):